# When F1 Fails: Granularity-Aware Evaluation for Dialogue Topic Segmentation

### v2

@@ 6.2 Boundary density (BOR)
Boundary Over-segmentation Ratio (BOR) measures the number of predicted boundaries relative to the number of gold boundaries. Although termed the boundary oversegmentation ratio, BOR is used here as a general boundary density ratio: values greater than 1 indicate oversegmentation relative to gold, while values less than 1 indicate undersegmentation. Formally, BOR is defined as the ratio $|P_{d}|/|G_{d}|$, computed per dialogue and aggregated across dialogues.
@@ 6.5 Baselines
We consider three classes of baselines: (i) a no-boundary baseline, (ii) periodic segmentation baselines, and (iii) oracle-density random baselines. Oracle-density baselines explicitly use gold boundary counts and are therefore not deployable; they are included solely to expose the sensitivity of boundary metrics to density alignment.
@@
Oracle-periodic Every-$N$ DialSeg711 ($N=4$) 0.639 0.97 0.812 0.823
@@
Table 3: Baseline diagnostic statistics relative to gold annotations. Oracle-density baselines match gold boundary counts by construction.
@@ 7 Results
This section evaluates the proposed separation between boundary scoring and boundary selection under differing annotation granularity. The results below instantiate multiple failure modes summarized in Table 2, including for previously published methods re-evaluated under the same framework.
@@ 7.2 External Methods Under Granularity-Aware Evaluation
To assess whether the granularity effects identified in this work are specific to our scoring model or reflect broader evaluation behavior, we re-evaluated a representative subset of published dialogue segmentation methods under the proposed granularity-aware framework. The selected methods include simple baselines (random and periodic segmentation), a classical unsupervised approach (TextTiling), and a coherence-based model (CSM). All methods were evaluated using the same boundary canonicalization and metrics as in previous sections.
@@
All single-point results reported below correspond to individual operating points along the density–quality curves shown in Figure 1.
@@
TextTiling DialSeg711 0.648 1.88 0.550 0.968 0.746
CSM DialSeg711 0.412 0.64 0.373 0.749 0.921
@@
Oracle-Periodic DialSeg711 0.483 1.00 0.211 0.806 0.809
TextTiling TIAGE 0.612 2.04 0.445 0.949 0.693
CSM TIAGE 0.615 3.97 0.346 0.976 0.454
@@
Oracle-Periodic TIAGE 0.620 1.00 0.300 0.832 0.838
Table 4: External dialogue segmentation methods re-evaluated under the proposed granularity-aware framework. Boundary density (BOR) varies substantially across methods and datasets, producing distinct density regimes despite comparable boundary accuracy scores.
Several patterns in Table 4 align with the failure modes summarized in Table 2. Simple periodic segmentation achieves competitive boundary accuracy when its output density aligns with the gold annotation (BOR $\approx 1$), despite lacking semantic grounding. Classical methods such as TextTiling exhibit recall-dominated behavior with elevated BOR, yielding high purity alongside aggressive oversegmentation. Notably, the same method (CSM) occupies different density regimes across datasets, indicating that apparent performance differences often reflect density mismatch rather than improved boundary detection.
@@
Table 5: Density regimes under the proposed failure taxonomy, derived from BOR thresholds.
@@
TT vs. CSM TIAGE +0.128 +0.219 [0.155, 0.283] +1.26 [1.11, 1.44] Aggressive $\rightarrow$ More Aggressive
@@
Table 6: Pairwise audit of runnable SuperDialseg leaderboard comparisons with 95% dialogue-level bootstrap confidence intervals. $T$ denotes TextTiling and Oracle-P is the Oracle-Periodic (Every-NN) baseline. Positive $\Delta$ indicates higher values for the first-listed method. Arrows denote relative changes in boundary density from the second-listed method to the first-listed method. Regime labels follow Table 2.
@@ 7.3 Stage 1: Synthetic Pretraining on Splice Boundaries
While some synthetic splice points may join semantically related segments, this introduces conservative label noise that discourages hallucinated boundaries rather than inducing spurious segmentation.
@@ 7.4 Stage 2: Supervised Fine-Tuning on Benchmarks
Epoch Overall F1 W-F1 BOR SuperSeg F1/BOR TIAGE F1/BOR
@@
Table 8: Stage 2 supervised fine-tuning on benchmark datasets. Boundary density converges toward annotated density (BOR $\approx 1$) on datasets seen during training, while W-F1 remains high.
@@ 7.5 Stage 3: Final Test with Calibration
Table 9: Stage 3 final test results. Boundary density (BOR) is reported alongside boundary accuracy to distinguish granularity mismatch from detection failure.
@@ 9 Discussion
Across datasets, we observe two stable annotation regimes: sparse gold (few annotated boundaries per dialogue) and dense gold (many annotated boundaries per dialogue).
When gold annotations are relatively uniform (e.g., DialSeg711), granularity errors concentrate in a small subset of atypical dialogues. In this setting, emitting additional boundaries is often beneficial, and W-F1 continues to improve under moderate oversegmentation.
@@
To confirm that granularity effects are not specific to our scorer, we evaluated CSM xing2021csm, a published coherence-based segmentation model, under the same framework. CSM operates in the conservative regime (BOR = 0.64), producing fewer boundaries than the gold annotation. Its low strict F1 (0.373) reflects low recall rather than incoherent segmentation: coverage remains high (0.921), indicating that predicted segments are valid but coarser than the reference. This demonstrates that the granularity–F1 interaction identified here generalizes across structurally distinct segmentation methods.
@@
These guidelines reinforce the central claim of this paper: boundary density is not an intrinsic property of a conversation, but a design choice that should be tuned to application requirements rather than optimized to match a given annotation scheme. This dependence on selection behavior motivates treating boundary selection as a distinct, explicit step.
@@ 9.3 On Dataset Choice and External Validity
Few large-scale corpora of human–LLM interaction are publicly available, and none currently include turn-level topic boundary annotations suitable for segmentation evaluation. WildChat zhao2024wildchat, for example, contains over one million authentic user–assistant exchanges but is released without discourse- or topic-level segmentation labels. Nevertheless, such data exhibit strong initiative asymmetries: users introduce goals and direct topic shifts, while assistant turns are reactive responses constrained by the preceding query. This asymmetry is well established in studies of institutional and task-oriented interaction drew1992talk and has been shown to shape discourse segmentation under mixed-initiative conditions acl1990mixedinitiative. In deployed AI assistant systems, these asymmetries are typically addressed through systems-level design choices for intent tracking and context management chopra2024exploring, which operate downstream of the boundary detection problem studied here.
@@ 10 Limitations and Future Work
We report dialogue-level bootstrap confidence intervals for selected figures and tables (e.g., Figure 1 and Tables 34) to characterize variability arising from dialogue sampling under fixed model checkpoints and boundary-selection parameters. These intervals are computed by resampling dialogues and do not reflect variability across random seeds, training runs, or alternative model initializations.
@@
Inter-annotator agreement (IAA) statistics could also provide additional empirical support for the claim that topic boundaries are granularity-dependent rather than objective. Several dialogue segmentation datasets report IAA in their original publications (e.g., DialSeg711 dialseg711, SuperSeg superdialseg2023), often showing only moderate agreement even under controlled annotation protocols. However, these statistics are reported using heterogeneous definitions and units (e.g., boundary placement, segment overlap, or task-level agreement), making direct comparison across datasets difficult. Because the present work focuses on evaluation behavior under fixed annotation schemes rather than on the reliability of those schemes themselves, we do not reanalyze annotator agreement here. Integrating IAA into a unified analysis of granularity effects remains an important direction for future work.
@@ 11 Conclusion
A central empirical finding of this work is that boundary-based metrics can be dominated by boundary density alignment rather than boundary placement quality. For example, we observe that the same published method (CSM) occupies different density regimes across datasets, a pattern consistent with differences in annotation density rather than changes in detection capability. Targeted audits of runnable SuperDialseg leaderboard comparisons further show that apparent gains in boundary accuracy frequently coincide with shifts toward more aggressive density regimes, rather than improved boundary localization at a matched granularity.
@@ Appendix D Boundary canonicalization
We provide per-dataset extraction details (file paths, fields, and filtering) in Table 11.
@@
Table 11: Boundary canonicalization details required to reproduce gold boundary sets $G_{d}$ from each dataset. Unit codes: U = utterance-level boundaries between consecutive turns from either speaker; T = speaker-turn boundaries.
@@ Appendix F Adaptive Boundary Selection
This implements a persistence test: a candidate’s evidence increases only if the same boundary signal remains present as more context is observed.
@@
The adaptive controller operates on candidate positions rather than messages, making the target rate invariant to dialogue length. Evidence accumulation uses a frozen score to test persistence of a boundary signal rather than repeatedly re-scoring shifting contexts. When $\tau(t)$ is held fixed and accumulation is disabled, the procedure reduces to the static selection rule used in the main experiments.

### v3

@@ 6.2 Boundary density (BOR)
Boundary Over-segmentation Ratio (BOR) measures the number of predicted boundaries relative to the number of gold boundaries. Although termed the boundary oversegmentation ratio, BOR is used here as a general boundary density ratio: values greater than 1 indicate oversegmentation relative to gold, while values less than 1 indicate undersegmentation. Formally, BOR is defined as the ratio of total predicted boundaries to total gold boundaries across all dialogues: $\mathrm{BOR}=\sum_{d}|P_{d}|\;/\;\sum_{d}|G_{d}|$. This micro-aggregation reflects overall boundary density rather than being skewed by high-variance ratios in short dialogues.
@@ 6.5 Baselines
We consider three classes of baselines: (i) a no-boundary baseline, (ii) periodic segmentation baselines (Periodic Every-$N$), which place boundaries at fixed intervals with $N$ chosen per-dataset, and (iii) oracle-density random baselines. We additionally evaluate an Oracle-Periodic baseline that matches the gold boundary count per dialogue exactly (used in Tables 5 and 7).
@@
Periodic Every-$N$ DialSeg711 ($N=4$) 0.639 0.97 0.812 0.823
@@
Table 4: Baseline diagnostic statistics relative to gold annotations. Periodic (Every-NN) places boundaries at fixed intervals, with $N$ chosen per-dataset to approximate gold density. Oracle-density baselines match gold boundary counts per dialogue by construction. Where non-zero, W-F1 for no-boundary reflects dialogues lacking gold boundaries (W-F1=1 under empty-set handling; see Appendix A).
@@ 7 Results
This section evaluates the proposed separation between boundary scoring and boundary selection under differing annotation granularity. The results below instantiate multiple failure modes summarized in Table 3, including results for previously published methods re-evaluated under the same framework.
@@ 7.2 External Methods Under Granularity-Aware Evaluation
To assess whether the granularity effects identified in this work are specific to our scoring model or reflect broader evaluation behavior, we re-evaluated a representative subset of published dialogue segmentation methods under the proposed granularity-aware framework. The selected methods include simple baselines (random and periodic segmentation), a classical unsupervised approach (TextTiling), and a coherence scoring model (CSM; xing2021improving), which uses BERT’s Next Sentence Prediction (NSP) task to score coherence between adjacent utterances. All methods were evaluated using the same boundary canonicalization and metrics as in previous sections.
@@
All single-point results reported below correspond to individual operating points along density–quality curves (see Figure 1 for DialSeg711; TIAGE curves follow the same pattern).
@@
TextTiling DialSeg711 0.613 2.69 0.488 0.988 0.638
CSM DialSeg711 0.393 0.56 0.353 0.732 0.931
@@
Oracle-Periodic DialSeg711 0.481 1.00 0.211 0.806 0.809
TextTiling TIAGE 0.628 1.84 0.434 0.933 0.717
CSM TIAGE 0.409 0.58 0.306 0.760 0.909
@@
Oracle-Periodic TIAGE 0.633 1.00 0.276 0.824 0.826
Table 5: External dialogue segmentation methods re-evaluated under the proposed granularity-aware framework. Oracle-Periodic places boundaries at evenly spaced intervals with count matched to gold per dialogue (BOR=1.00 by construction). Boundary density (BOR) varies substantially across methods, producing distinct density regimes despite comparable boundary accuracy scores. All threshold-based methods (TextTiling, CSM) evaluated at fixed operating points; no per-dataset tuning.
Several patterns in Table 5 align with the failure modes summarized in Table 3. Simple periodic segmentation achieves competitive boundary accuracy when its output density aligns with the gold annotation (BOR $\approx 1$), despite lacking semantic grounding. Classical methods such as TextTiling exhibit recall-dominated behavior with elevated BOR, yielding high purity alongside aggressive oversegmentation. CSM operates conservatively on both datasets (BOR $\approx 0.56$$0.58$), producing high coverage but lower purity—a precision-dominated pattern characteristic of undersegmentation.
@@
Table 6: Density regimes under the proposed failure taxonomy, derived from BOR thresholds. Regime classifications based on fixed operating points (no per-dataset tuning).
Threshold tuning and density side effects
The results in Tables 5 and 6 use fixed internal thresholds for TextTiling and CSM, isolating intrinsic method behavior from threshold effects. Standard practice, however, tunes thresholds on a development set to maximize F1. We find that such tuning can produce dramatically different operating points across datasets. For example, when CSM’s threshold is tuned per-dataset, the method shifts from conservative (BOR =0.64=0.64) on DialSeg711 to aggressive (BOR =3.97=3.97) on TIAGE—a sixfold increase in boundary density. With a fixed threshold, CSM operates conservatively on both datasets (BOR $\approx 0.56$$0.58$), revealing that the regime variation is a tuning artifact rather than inherent method behavior. F1-based tuning selects boundary density as a side effect; the resulting operating point can vary substantially across datasets depending on annotation density, which is precisely the confound that BOR-aware evaluation is designed to expose.
@@
TT vs. CSM TIAGE +0.128 +0.219 [0.155, 0.283] +1.26 [1.11, 1.44] Conservative $\rightarrow$ Aggressive
@@
Table 7: Pairwise audit of runnable SuperDialseg leaderboard comparisons with 95% dialogue-level bootstrap confidence intervals. $T$ denotes TextTiling and Oracle-P is the Oracle-Periodic baseline (boundary count matched to gold per dialogue). Positive $\Delta$ indicates higher values for the first-listed method. Arrows denote relative changes in boundary density from the second-listed method to the first-listed method. Regime labels follow Table 6. All methods evaluated at fixed operating points.
@@ 7.4 Stage 2: Supervised Fine-Tuning on Benchmarks
Epoch Overall (both datasets) SuperSeg TIAGE
F1 W-F1 BOR F1 / BOR F1 / BOR
@@
Table 9: Stage 2 supervised fine-tuning on benchmark datasets. Metrics are computed on the held-out validation split (not used for gradient updates). “Overall” F1 and BOR are micro-averaged across both datasets; W-F1 is macro-averaged over dialogues (Appendix A). Boundary density converges toward annotated density (BOR $\approx 1$) on datasets seen during training, while W-F1 remains high.
@@ 7.5 Stage 3: Final Test with Calibration
Table 10: Stage 3 final test results. Boundary density (BOR) is reported alongside boundary accuracy to distinguish granularity mismatch from detection failure. (F1 is micro-averaged; W-F1 is macro-averaged. See Appendix A.)
@@ 9 Discussion
The strength of this density–metric coupling varies across datasets depending on annotation consistency. When gold annotations are relatively uniform across dialogues (e.g., DialSeg711), granularity errors concentrate in a small subset of atypical dialogues. In this setting, emitting additional boundaries is often beneficial, and W-F1 continues to improve under moderate oversegmentation.
@@
These guidelines reinforce the central claim of this paper: boundary density is not an intrinsic property of a conversation, but a design choice that should be tuned to application requirements rather than optimized to match a given annotation scheme. As Section 7 demonstrates, F1-based threshold tuning implicitly searches for the density that matches gold annotations rather than improving boundary localization. This dependence on selection behavior motivates treating boundary selection as a distinct, explicit step.
@@ 9.3 On Dataset Choice and External Validity
Few large-scale corpora of human–LLM interaction are publicly available, and none currently include turn-level topic boundary annotations suitable for segmentation evaluation. WildChat zhao2024wildchat, for example, contains over one million authentic user–assistant exchanges but is released without discourse- or topic-level segmentation labels. Nevertheless, such data exhibit strong initiative asymmetries: users introduce goals and direct topic shifts, while assistant turns are reactive responses constrained by the preceding query. This asymmetry is well established in studies of institutional and task-oriented interaction drew1992talk and has been shown to shape discourse segmentation under mixed-initiative conditions walker1990mixed. In deployed AI assistant systems, these asymmetries are typically addressed through systems-level design choices for intent tracking and context management chopra2024exploring, which operate downstream of the boundary detection problem studied here.
@@ 10 Limitations and Future Work
We report dialogue-level bootstrap confidence intervals for selected figures and tables (e.g., Figure 1 and Table 7) to characterize variability arising from dialogue sampling under fixed model checkpoints and boundary-selection parameters. These intervals are computed by resampling dialogues and do not reflect variability across random seeds, training runs, or alternative model initializations.
@@
Inter-annotator agreement (IAA) statistics could also provide additional empirical support for the claim that topic boundaries are granularity-dependent rather than objective. Several dialogue segmentation datasets report IAA in their original publications (e.g., DialSeg711 xu2021topicaware, SuperSeg jiang2023superdialseg), often showing only moderate agreement even under controlled annotation protocols. However, these statistics are reported using heterogeneous definitions and units (e.g., boundary placement, segment overlap, or task-level agreement), making direct comparison across datasets difficult. Because the present work focuses on evaluation behavior under fixed annotation schemes rather than on the reliability of those schemes themselves, we do not reanalyze annotator agreement here. Integrating IAA into a unified analysis of granularity effects remains an important direction for future work.
@@ 11 Conclusion
A central empirical finding of this work is that boundary-based metrics can be dominated by boundary density alignment rather than boundary placement quality. For example, under F1-tuned operating points, CSM occupied markedly different density regimes across datasets (conservative on DialSeg711, aggressive on TIAGE); under fixed operating points, its boundary density was stable. The observed regime shift was a tuning artifact, not an intrinsic method–dataset interaction—illustrating how F1-based optimization can actively induce the granularity confounds that standard evaluation protocols fail to diagnose.
@@ Appendix D Boundary canonicalization
Table 12 summarizes the canonicalization rules; complete extraction implementations are in the accompanying code repository.
@@
Table 12: Boundary canonicalization details required to reproduce gold boundary sets $G_{d}$ from each dataset. Unit codes: U = each message is a boundary candidate (typical for two-party dialogues where speakers alternate); T = boundaries occur only at speaker changes, grouping consecutive sentences from the same speaker into one unit (used for multi-party meeting transcripts like QMSum).
@@ Appendix F Adaptive Boundary Selection
This implements a time-to-threshold integrator: candidates with higher initial scores $s_{i}$ reach the commitment threshold faster than candidates with lower scores. Freezing the score at the time of first availability ensures that the commitment decision reflects the strength of the boundary signal at the point of topic departure, rather than being confounded by subsequent context that may diverge further or return to the original topic. The adaptive threshold $\tau(t)$ (below) provides the primary density control mechanism.
@@
The adaptive controller operates on candidate positions rather than messages, making the target rate invariant to dialogue length. Freezing the boundary score avoids repeatedly re-scoring shifting contexts, and ensures that high-scoring candidates are committed before low-scoring ones under time pressure imposed by the adaptive threshold. When $\tau(t)$ is held fixed and accumulation is disabled, the procedure reduces to the static selection rule used in the main experiments.